Calculate the probabilities of P(-3.10<z<1.90)
1
Expert's answer
2021-09-29T08:46:39-0400

"P(-3.10<z<1.90)= P(0<z<1.9)+P(-3.1<z<0)"

Required probabilities can be found in table of z distribution

"P(-3.1<z<0) = 0.49903"

"P(0<z<1.9) = 0.47128"

"P(-3.10<z<1.90) = 0.49903+0.47128=0.97031"